# Can Large Language Models Develop Gambling Addiction?

### v1

@@ E.2 LLM Sparse Autoencoder
Recently, the Sparse Autoencoder (SAE) technique has been rapidly emerging as a core tool in LLM interpretability research. Cunningham et al. (2024) showed that by applying SAE to the internal activation values (residual stream) of LLMs, it is possible to resolve the problem of polysemanticity (where a single neuron represents a mix of multiple semantic functions), which was a problem in existing neural networks. By enforcing sparse activations through regularization, SAE finds interpretable directions where one feature has one clear meaning (a monosemantic feature). In particular, it showed superior results in automated interpretability scores compared to existing methods (PCA, ICA, etc.), and experimentally proved that it can finely specify which activation features play a causal role in downstream tasks, such as identifying indirect objects within actual sentences. As such, SAE-based decomposition has a distinct significance in that it effectively solves the problem of superposition within LLMs and is scalable with only large-scale unsupervised data.
@@
Shi et al. (2025) overcame the limitation of existing SAEs that extract features from only a single layer and proposed a “routing” structure that integrates and weights activation information across multiple layers. This dramatically improved feature interpretability and enabled the analysis of semantic flows and interactions between layers. Meanwhile, Anthropic’s Sparse Crosscoder (Lindsey et al., 2024) and OpenAI’s Scaling SAE (Adamek et al., 2025) are contributing to the practical improvement of model transparency and reliability by intensively supplementing aspects such as training stability, feature deduplication, and evaluation metric improvements required for applying SAE to entire large-scale LLMs.
@@
Appendix F LLM Usage
We utilized Large Language Models (LLMs) to support various aspects of this research. Specifically, we employed Anthropic’s Claude (Anthropic, 2025) for surveying previous research, assisting with code implementation, cleaning data, and generating figures from the processed data. For improving the grammar and clarity of expression in the manuscript, we used Google’s Gemini (Gemini Team and Google, 2025). The authors have reviewed and taken full responsibility for all content, including any text or code generated with the assistance of these models.

### v2

@@ D.2 LLM Sparse Autoencoder
Recently, the Sparse Autoencoder (SAE) technique has been rapidly emerging as a core tool in LLM interpretability research. cunningham2024sparse showed that by applying SAE to the internal activation values (residual stream) of LLMs, it is possible to resolve the problem of polysemanticity (where a single neuron represents a mix of multiple semantic functions), which was a problem in existing neural networks. By enforcing sparse activations through regularization, SAE finds interpretable directions where one feature has one clear meaning (a monosemantic feature). In particular, it showed superior results in automated interpretability scores compared to existing methods (PCA, ICA, etc.), and experimentally proved that it can finely specify which activation features play a causal role in downstream tasks, such as identifying indirect objects within actual sentences. As such, SAE-based decomposition has a distinct significance in that it effectively solves the problem of superposition within LLMs and is scalable with only large-scale unsupervised data.
@@
shi2025route overcame the limitation of existing SAEs that extract features from only a single layer and proposed a “routing” structure that integrates and weights activation information across multiple layers. This dramatically improved feature interpretability and enabled the analysis of semantic flows and interactions between layers. Meanwhile, Anthropic’s Sparse Crosscoder (lindsey2024crosscoders) and OpenAI’s Scaling SAE (adamek2024scaling) are contributing to the practical improvement of model transparency and reliability by intensively supplementing aspects such as training stability, feature deduplication, and evaluation metric improvements required for applying SAE to entire large-scale LLMs.
@@
Appendix E LLM Usage
We utilized Large Language Models (LLMs) to support various aspects of this research. Specifically, we employed Anthropic’s Claude (anthropic2025claude4) for surveying previous research, assisting with code implementation, cleaning data, and generating figures from the processed data. For improving the grammar and clarity of expression in the manuscript, we used Google’s Gemini (google2025gemini2_5). The authors have reviewed and taken full responsibility for all content, including any text or code generated with the assistance of these models.